PET has a Young's, Modulus of 3.5 GPa and PLGA 1.2 GPa. A polymer blend is prepared with 50:50 mixture (PET:PlGA) . What is the Young's, Modulus of this blend (use mixture rule)? Calculate the percent load borne by PET when the blend is subjected to a load. If after 10 days, half of PET and one third of PLGA had degraded, what is the Young's Modulus of this blend ?
1
Expert's answer
2020-04-17T10:48:59-0400

According to mixture rule, the Young's modulus of this blend will be


"Y=Y_{PET}\cdot0.5+Y_{PLGA}\cdot0.5=2.36\text{ GPa}."

The percent load borne by PET when the blend is subjected to a load is


"\epsilon_\%=\frac{Y_\text{PET}\cdot0.5}{Y}\cdot100\%=\\ \space\\=\frac{3.5\cdot0.5}{2.36}\cdot100\%=74\%."

If after 10 days, half of PET and one third of PLGA had degraded, the Young's Modulus of this blend will be


"Y_{10}=\\=Y_\text{PET}\cdot0.5\cdot(1-0.5)+Y_\text{PLGA}\cdot0.5\cdot(1-1/3)=\\=1.28\text{ GPa}."